OpenStudy (ammarah):

Write a rule for the nth term of the geometric sequence. Find a6. Then grapth the first five terms of the sequences r=3, a1=2

OpenStudy (ammarah):

@douglaswinslowcooper

OpenStudy (anonymous):

\[a,ar,ar^2,ar^3,ar^4,...\] \[2,2\times 3,2\times 3^2,2\times 3^3,2\times 3^4,...\] is what this looks like general term is \(a_n=2\times 3^{n-1}\)

OpenStudy (ammarah):

?

OpenStudy (ammarah):

@ganeshie8

OpenStudy (anonymous):

As @satellite73 indicated , the first term will be a1 = 2, and the next ones will be a2=(a1)(3) = 6, a3=(a2)(3)= (a1)(3)^2= (2)(9) =18 an = (a1)(3)^(n-1) = (2)(3)^(n-1)

OpenStudy (ammarah):

so thats the answer? how about the graphing part

OpenStudy (primeralph):

Do you know how to plot graphs?